LASER INDUCED FLUORESCENCE STUDIES OF JET-COOLED CF_{2}: DETERMINATION OF Ã-STATE STRETCHING FREQUENCIES

\underline{CHUJI WANG}, CONGXIANG CHEN, JINGHUA DAI, and XINGXIAO MA, Department of Chemical Physics, University of Science and Technology of China, Hefei, Anhui 230026, P. R. China.

Laser induced fluorescence (LIF) excitation spectra of CF_{2}Ã^{1}B_{1} ← X̃^{1}A_{1} system around 250 nm has been measured under supersonic free-jet conditions. CF_{2} radical was produced by DC-discharge of CF_{4} seeded in argon. The spectra present the overall spectral features of this system, from the band origin to Ã -state predissociation threshold that is measured to be 5.4eV. The previously ambiguous assignments of Ã-state stretching frequencies have been experimentally determined to be ν′_{1} = 1012.1 cm^{−1}, ν′_{3} = 1180.2 cm^{−1} , by analyzing the rotational subband spectra attributed to some of the observed progressions (2_{0}^{n}, 1_{0}^{1}2_{0}^{n−2} and 2_{0}^{n}3_{1}^{1} (n ≤ 11)) in which 30 vibronic bands were first unambiguously labeled. The measured Ã-state stretching frequencies do not support King’s [J. Mol. Spectry, 1, 78(1979)] tentative assignments, but are in good agreement with Cameron’s [J. Chem. Phys. 4475, 103(1995)] theoretical assertions. Furthermore, the systematic and detail measurements of the rotationally resolved LIF spectrum of this system for all the transitions in 2_{0}^{n} and 1_{0}^{1}2_{0}^{n−2} progressions have been conducted. The rotational and centrifugal constants of all (0, n, 0) and (1, n − 2, 0) vibrational levels in Ã-state are derived by least-squares fit. Some negative inertial defects are tentatively explained resulted from local perturbation.